OpenStudy (anonymous):

Find a quadratic function in standard form that has the following points: (1, -2); (2, -2); (3, -4).

OpenStudy (tkhunny):

Can you find the Axis of Symmetry without doing any algebra?

OpenStudy (anonymous):

im not sure

OpenStudy (tkhunny):

Shall we guess that (0,-4) is yet another point?

OpenStudy (tkhunny):

Axis of Symmetry is x = 3/2. Please tell me why?

OpenStudy (tkhunny):

Quadratic Function in Standard Form: \(f(x) = ax^{2} + bx + c\) From your three points: (1, -2); (2, -2); (3, -4). We have: \(f(x) = a(1)^{2} + b(1) + c = -2\) \(f(x) = a(2)^{2} + b(2) + c = -2\) \(f(x) = a(3)^{2} + b(3) + c = -4\) And you have some algebra in your future!